3. A division wide aptitude test in Mathematics was conducted to 3 800 students. The mean of the test is 58 and the standard deviation is 14.



a. What is the score that divides the distribution into two such that 80% of the cases is below it.



b. What are the scores that bounds the middle 90% of the distribution.



c. How many scores are between 60 and 96?



d. How many students belong to the top 12% of the examinees?






1
Expert's answer
2022-04-27T10:51:05-0400

Let "X=" score, "X\sim N(\mu, \sigma^2)."

a.

"P(X<x)=P(Z<\dfrac{x-\mu}{\sigma})"

"=P(Z<\dfrac{x-58}{14})=0.8"

"\dfrac{x-58}{14}\approx0.841621"

"x=69.78"

b.


"P(X<x_1)=P(Z<\dfrac{x_1-\mu}{\sigma})"

"=P(Z<\dfrac{x_1-58}{14})=0.05"

"\dfrac{x_1-58}{14}\approx-1.6449"

"x_1=35"


"P(X>x_2)=P(Z>\dfrac{x_2-\mu}{\sigma})"

"=P(Z>\dfrac{x_2-58}{14})=0.05"

"\dfrac{x_2-58}{14}\approx1.6449"

"x_2=81"

Between 35 and 81.


c.


"P(60<X<96)"

"=P(Z<\dfrac{96-58}{14})-P(Z\le\dfrac{60-58}{14})"

"\approx0.99668-0.55680\approx0.43988"

"0.43988(3800)=1672"

1672 scores are between 60 and 96.


d.


"0.12(3800)=456"

456 students belong to the top 12% of the examinees.